OpenStudy (anonymous):

A 250 g object hangs from a spring and oscillates with an amplitude of 5.42 cm. The spring constant is 48.0 N/m. b. What is the maximum speed of the object? c. At what position will the maximum speed occur?

OpenStudy (festinger):

Amplitude is the maximum displacement from the equilbrium position. System is a spring, assume it obey's hookes law: \[F=-kx\]and \[U_{elastic}=\frac{1}{2}kx^{2}\] At any point the energy is: \[\frac{1}{2}kx^{2}+\frac{1}{2}mv^{2}= Constant\] By conservation of energy, \[U_{max}=KE_{max}\] This occurs when U is 0, and this occurs when x=0 (can you see why?) Thus to find maximum speed simple equate (after cancelling the 1/2) \[kA^{2}=mv^{2}\] where A is the amplitude.

OpenStudy (anonymous):

You can use \[E \max = 1/2 mw ^{2}A ^{2} = I/2 m V ^{2}\], SO You have \[V = w A\] also from F=\[-kx =m a\] i.e\[w ^{2}=k/m\] \[w=\sqrt{k/m}\]=\[\sqrt{192}=13.9 s ^{-1}\] so, V max =13.9 * 5.42=75.338 m/s----occurs at the mean position i.e at unstratched position